ADNAN MUHAMMED

Grade 12,

Which of the following pairs represent the same rational number? (-2/-3) and (2/3)

Which of the following pairs represent the same rational number? 
 (-2/-3) and (2/3)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to check the given pair represents the same rational number.
​Then,
-2/-3 = 2/3
2/3= 2/3
∵2/3 = 2/3
∴-2/-3 = 2/3
So, the given pair is represents the same rational number.

Thanks